### v1

# Cardinality-Preserving Structured Sparse Graph Transformers for Molecular Property Prediction

Abstract
Drug discovery motivates efficient molecular property prediction under limited labeled data. Chemical space is vast (often estimated at $\sim 10^{60}$ drug-like molecules) while only thousands of drugs have been approved [6, 1]. Self-supervised pretraining on large unlabeled corpora has become essential for data-efficient molecular representation learning [10, 3]. We introduce CardinalGraphFormer, a graph transformer that adopts Graphormer-inspired structural biases (SPD and centrality; direct-bond edge bias) within a structured sparse attention regime (SPD $\leq 3$) and augments it with a cardinality-preserving unnormalized aggregation channel over the same support set. Pretraining combines contrastive graph-level alignment with masked attribute reconstruction. Under a fully matched evaluation protocol, CardinalGraphFormer improves mean performance on all 11 tasks and achieves statistically significant gains on 10 of 11 public benchmarks spanning MoleculeNet, OGB, and TDC ADMET tasks over strong reproduced baselines [8, 4, 5].
1 Introduction
Molecular property prediction serves as a key proxy task in drug discovery, constrained by limited experimental labels and an immense candidate space [6, 1]. Graph transformers with self-supervised pretraining have shown strong performance on standardized benchmarks [9, 10, 3].
We adopt Graphormer-inspired structural biases (SPD + centrality; direct-bond edge bias) but implement structured sparse attention using $K$-hop SPD masks for locality and efficiency. We further augment this with a cardinality-preserving channel to retain dynamic support-size signals complementary to static centrality embeddings.
Contributions
  • A structured sparse graph transformer layer with aligned cardinality-preserving attention.
  • Dual-objective self-supervised pretraining on $\sim$28M molecules.
  • Rigorous matched-protocol evaluation showing consistent, often significant gains over reproduced baselines. Code and artifacts are provided in the repository; a stable release commit will be frozen upon acceptance.
2 Related Work
Message passing remains a dominant paradigm for molecular graphs [2], and attention-based GNNs are widely used [7]. Graph transformers such as Graphormer inject structural biases (shortest-path distance, edge features, centrality) into global attention [9]. Sparse variants often restrict attention to local neighborhoods (e.g., $K$-hop masking) to reduce computation in practice. Cardinality-preserving attention adds unnormalized channels to retain degree/cardinality signals [11]. Self-supervised pretraining on graphs includes contrastive methods (GraphCL) and masked modeling (GraphMAE) [10, 3]. CardinalGraphFormer unifies structured sparse attention with aligned cardinality preservation and dual-objective pretraining.
3 Methods
3.1 Notation
[…]
3.2 Structured Sparse Attention with Graphormer-Inspired Biases
The support set is $\mathcal{S}(i)=\{j\mid\mathrm{SPD}(i,j)\leq K\}$ (includes self, SPD(i,i)=0(i,i)=0). Main experiments use $K=3$. Attention is masked to $\mathcal{S}(i)$ (outside: -$\infty$). Logits:
$a_{ij}=\frac{q_{i}^{\top}k_{j}}{\sqrt{d_{h}}}+b_{\phi(\mathrm{SPD}(i,j))}+b_{e_{ij}}+b_{c}(j),$ (1)
where $\phi(d)=\min(d,K)$ yields $K+1$ bins $\{0,\dots,K\}$, $b_{e_{ij}}$ is a learnable embedding of direct-bond features (bond type, aromaticity, conjugation, ring membership, and stereo configuration; categorical fields are one-hot encoded and embedded separately, then summed; stereo uses an explicit ‘none’ category when unspecified or absent; edge features are symmetric for the undirected graph; $b_{e_{ij}}=0$ for SPD>1>1), and $b_{c}(j)$ is key-side degree-bin embedding. Complexity scales as $O(N\cdot\overline{|\mathcal{S}(i)|})$ per layer.
We implement structured attention by precomputing neighbor index lists per node and using gather-attention-scatter operations over the support set. To remain compatible with optimized attention kernels (e.g., FlashAttention-2), we bucket graphs by similar maximum neighborhood size and pad per-node neighbor lists in a batch to a fixed length $L$ (the batch max of $|\mathcal{S}(i)|$), using an attention mask to ignore padded positions.
3.3 Cardinality-Preserving Attention
Per-head output (layer $\ell$, head $m$):
$o_{i}^{(\ell,m)}=\sum_{j\in\mathcal{S}(i)}\alpha_{ij}^{(\ell,m)}v_{j}^{(\ell,m)}+g_{i}^{(\ell,m)}\odot\sum_{j\in\mathcal{S}(i)}v_{j}^{(\ell,m)},$ (2)
where $\alpha_{ij}^{(\ell,m)}=\frac{\exp(a_{ij}^{(\ell,m)})}{\sum_{k\in\mathcal{S}(i)}\exp(a_{ik}^{(\ell,m)})}$ and both sums use the head-specific value projections $v_{j}^{(\ell,m)}$ (including self). Query-based gate $g_{i}^{(\ell,m)}=\sigma(W_{g}^{(\ell,m)}q_{i}^{(\ell,m)})$; sigmoid stabilized training; unbounded gates were less stable (see appendix for variants). Concatenation and projection yield $u_{i}^{(\ell)}$, then static centrality injection $\tilde{u}_{i}^{(\ell)}=u_{i}^{(\ell)}+c_{i}$ where $c_{i}=E_{c}(\mathrm{deg\_bin}(i))$. Canonical residual blocks follow (attention + FFN, each with Dropout + Add + LayerNorm). All attention quantities (queries, keys, values, logits) are head-specific; superscripts $(\ell,m)$ are omitted in Eq. 1 for readability.
The degree-normalized CPA ablation (Table 6) divides the unnormalized sum by $|\mathcal{S}(i)|$ before gating and underperforms, consistent with the hypothesis that raw magnitude variation—correlated with current support size—carries useful information beyond static degree bins or explicit normalization. Supporting this, the L2 norm of the gated unnormalized sum (i.e., $\|g_{i}^{(\ell,m)}\odot\sum_{j\in\mathcal{S}(i)}v_{j}^{(\ell,m)}\|_{2}$, averaged over heads, computed before dropout and output projection) correlates positively with $|\mathcal{S}(i)|$ even after controlling for node degree (partial correlation $r\approx 0.52$; partial Pearson via residualization on degree bin), computed on a held-out 50k-graph pretraining validation subset ($\approx$1.1M nodes), averaged across 5 seeds; similar correlations observed across layers 3–10.
3.4 Self-Supervised Objectives
Contrastive views. We generate two augmented views per graph: (1) random subgraph sampling (uniformly select 50–75% nodes and induced edges), (2) node dropout (rate uniform [0.1,0.3] on surviving nodes), (3) edge dropout (rate uniform [0.1,0.3] on surviving edges). SPD distances and degree bins are computed once on the original graph and preserved across views. For each view, $\mathcal{S}(i)$ is intersected with surviving nodes. This produces views that differ in attributes and connectivity while preserving a stable locality prior used by the model; recomputing SPD per view would be prohibitively expensive and introduce instability in the masking pattern. When edges are dropped, the attention mask remains based on original topology, effectively treating dropped edges as temporary perturbations within a fixed structural scaffold. This is a deliberate design choice that stabilizes training; ablations with per-view SPD recomputation showed similar or slightly worse downstream performance (see appendix).
Graph readout: mean-pool final node embeddings; 2-layer MLP projection; NT-Xent ($\tau=0.2$) with in-batch negatives.
Masked modeling. 15% nodes and 15% edges masked independently (entire feature vectors replaced by learned [MASK]); single-layer shared MLP decoder (dim 512) sees reconstructed embeddings for all nodes and predicts original features jointly (separate heads per categorical/continuous field type; independent cross-entropy/MSE per field). GraphMAE-K3 uses the same masking scheme and decoder architecture.
Total loss $\mathcal{L}=\mathcal{L}_{\text{mask}}+0.5\mathcal{L}_{\text{contrast}}$.
3.5 Model Architecture and Pretraining Details
12 layers, hidden 512, 8 heads, FFN 2048, LayerNorm, dropout 0.1. All hyperparameters fixed across models. Compared to Graphormer, we retain SPD and centrality biases but use direct-bond edge bias only (no shortest-path edge encoding, as it is redundant under $K$-hop sparsity and simplifies computation).
3.6 Pretraining Corpus and Standardization
We pretrain on $\sim$28M molecules (after filtering).
  • Source datasets: ZINC20 clean-leads (220M SMILES) and ChEMBL35 (2.4M molecules).
  • Licensing summary: ZINC20 (public domain); ChEMBL (CC-BY-SA 3.0).
  • Standardization: deterministic RDKit pipeline producing canonical isomeric SMILES; heavy atoms only.
  • Filtering/deduplication: removed invalid SMILES, salts/mixtures, non-drug-like (Lipinski violations); dedup by InChIKey; benchmark overlap removal.
3.7 Optimization and Training Schedule
  • Optimizer: AdamW.
  • Learning rate schedule: 2e-4 peak with 10k warmup steps, cosine decay to 1e-6.
  • Batch size and accumulation: 4096 graphs, no accumulation.
  • Epochs/steps: 500k pretraining steps; fine-tuning max 300 epochs with patience 50.
  • Regularization: weight decay 1e-5, dropout 0.1, no clipping.
  • Seeds: 5 fixed seeds (42, 43, 44, 45, 46) for reproducibility.
4 Experiments
All models share architecture, hyperparameters, pretraining schedule, corpus, and fine-tuning protocol. Differences are summarized in Table 1.
Model Pretraining Objective Attention Aggregation Support Mask ($K$) SPD / Bond Bias Key-side $b_{c}(j)$ Per-layer $c_{i}$
CardinalGraphFormer Mask + Contrast Normalized + CPA 3 Yes Yes Yes
SparseGraphormer-K3 Mask + Contrast Normalized only 3 Yes Yes Yes
GraphMAE-K3 Mask only Normalized only 3 Yes Yes Yes
Table 1: Model configurations. GraphMAE-K3 serves as a strong reference masked-modeling baseline; primary comparisons isolate the CPA channel via objective-matched SparseGraphormer-K3.
4.1 Baseline Parity
We match baselines to CardinalGraphFormer under a shared protocol.
  • Shared encoder/backbone: identical 12-layer transformer architecture and parameterization.
  • Pretraining data parity: same corpus and filtering.
  • Augmentations and objectives: GraphMAE uses the same masking; GraphCL uses the same contrastive pipeline but no masking.
  • Fine-tuning protocol: scaffold splits, early stopping on validation metric, same tuning grid.
4.2 Benchmarks and Metrics
We evaluate on 11 public molecular property tasks spanning MoleculeNet (ESOL, Lipophilicity, BBBP, Tox21, ClinTox) [8], OGB (ogbg-molhiv, ogbg-molpcba) [4], and TDC ADMET (Caco2_Wang, hERG, Clearance_Hepatocyte_AZ, Clearance_Microsome_AZ) [5]. Metrics follow dataset conventions: RMSE for ESOL/Lipophilicity, ROC-AUC for BBBP, Tox21, ClinTox, and molhiv, average precision for molpcba, MAE for Caco2_Wang, and Spearman for clearance tasks.
4.3 Efficiency
Setting Peak Memory (GB) Throughput (graphs/s) Mean NN Median NN Mean $|\mathcal{S}(i)|$ Median $|\mathcal{S}(i)|$ Median coverage (%)
$K=3$ (full corpus) 18.4 820 25 22 24 21 95
[…]
$K=3$ (large slice, $N\geq 50$) 22.1 610 68 62 52 49 79
[…]
Table 2: Efficiency measured on A100 40GB GPUs, bf16 precision, batch=4096 graphs, no activation checkpointing, FlashAttention-2 kernel. Both settings use the same attention implementation; $K=3$ is implemented as an attention mask (not block-sparse kernels); compute savings therefore come primarily from shorter effective attention spans on larger graphs and reduced quadratic scaling. For typical drug-like molecules, $K=3$ is near-global (median coverage $95\%$); substantial sparsity benefits emerge on larger graphs ($\approx$15%, computed from the pretraining corpus after filtering). Median coverage (%) is computed as $\mathrm{median}_{G}\!\left(\frac{1}{N}\sum_{i\in V}\frac{|\mathcal{S}(i)|}{N}\right)\times 100$.
Table 2 reports measured efficiency.
4.4 Results
Statistical significance uses paired bootstrap (10,000 resamples per task/seed, with identical resample indices across models to preserve pairing, deltas averaged across seeds); denotes 95% CI excludes zero vs SparseGraphormer-K3. Holm correction is applied across the 11 benchmark tasks for the primary CPA-vs-noCPA comparisons (conclusions unchanged).
Task (metric) SparseGraphormer-K3 $\Delta$ GraphMAE-K3 GraphCL CardinalGraphFormer
ESOL (RMSE $\downarrow$) 0.598 $\pm$ 0.03 -0.056 0.580 $\pm$ 0.02 0.612 $\pm$ 0.03 0.542 $\pm$ 0.02
Lipophilicity (RMSE $\downarrow$) 0.658 $\pm$ 0.02 -0.029 0.662 $\pm$ 0.02 0.672 $\pm$ 0.02 0.629 $\pm$ 0.01
BBBP (ROC-AUC $\uparrow$) 0.918 $\pm$ 0.011 +0.020 0.922 $\pm$ 0.010 0.910 $\pm$ 0.012 0.938 $\pm$ 0.009
Tox21 (ROC-AUC $\uparrow$) 0.908 $\pm$ 0.009 +0.013 0.909 $\pm$ 0.007 0.902 $\pm$ 0.010 0.921 $\pm$ 0.008
ClinTox (ROC-AUC $\uparrow$) 0.939 $\pm$ 0.017 +0.015 0.938 $\pm$ 0.015 0.932 $\pm$ 0.018 0.954 $\pm$ 0.011
Table 3: MoleculeNet scaffold split results (5 seeds). $\Delta=$ CardinalGraphFormer - SparseGraphormer-K3. For $\downarrow$ metrics, negative $\Delta$ indicates improvement. significant vs SparseGraphormer-K3.
Dataset (metric) SparseGraphormer-K3 $\Delta$ GraphMAE-K3 GraphCL CardinalGraphFormer
ogbg-molhiv (ROC-AUC $\uparrow$) 0.802 $\pm$ 0.009 +0.017 0.803 $\pm$ 0.008 0.798 $\pm$ 0.010 0.819 $\pm$ 0.007
ogbg-molpcba (AP $\uparrow$) 0.294 $\pm$ 0.002 +0.010 0.295 $\pm$ 0.002 0.292 $\pm$ 0.003 0.304 $\pm$ 0.002
Table 4: OGB official splits (5 seeds). $\Delta=$ CardinalGraphFormer - SparseGraphormer-K3. significant vs SparseGraphormer-K3.
Benchmark (metric) SparseGraphormer-K3 $\Delta$ GraphMAE-K3 GraphCL CardinalGraphFormer
Caco2_Wang (MAE $\downarrow$) 0.265 $\pm$ 0.009 -0.024 0.268 $\pm$ 0.008 0.270 $\pm$ 0.009 0.241 $\pm$ 0.007
hERG (ROC-AUC $\uparrow$) 0.870 $\pm$ 0.005 +0.028 0.872 $\pm$ 0.004 0.865 $\pm$ 0.006 0.898 $\pm$ 0.004
Clearance_Hepatocyte_AZ (Spearman $\uparrow$) 0.525 $\pm$ 0.020 +0.013 0.518 $\pm$ 0.018 0.520 $\pm$ 0.021 0.538 $\pm$ 0.017
Clearance_Microsome_AZ (Spearman $\uparrow$) 0.614 $\pm$ 0.012 +0.007 0.615 $\pm$ 0.012 0.610 $\pm$ 0.013 0.621 $\pm$ 0.011
Table 5: TDC ADMET scaffold split results (5 seeds). $\Delta=$ CardinalGraphFormer - SparseGraphormer-K3. For $\downarrow$ metrics, negative $\Delta$ indicates improvement. significant vs SparseGraphormer-K3.
4.5 Ablation Study
Variant molhiv (AUC) Caco2 (MAE) molpcba (AP)
CardinalGraphFormer (full, $K=3$) 0.819 $\pm$ 0.007 0.241 $\pm$ 0.007 0.304 $\pm$ 0.002
SparseGraphormer-K3 (no CPA) 0.802 $\pm$ 0.009 0.265 $\pm$ 0.009 0.294 $\pm$ 0.002
Degree-normalized CPA 0.810 $\pm$ 0.008 0.252 $\pm$ 0.008 0.298 $\pm$ 0.002
No per-layer $c_{i}$ 0.816 $\pm$ 0.007 0.245 $\pm$ 0.007 0.302 $\pm$ 0.002
No key-side $b_{c}(j)$ 0.817 $\pm$ 0.007 0.243 $\pm$ 0.007 0.303 $\pm$ 0.002
$K=2$ 0.812 $\pm$ 0.008 0.249 $\pm$ 0.008 0.300 $\pm$ 0.002
$K=5$ 0.816 $\pm$ 0.007 0.244 $\pm$ 0.007 0.302 $\pm$ 0.002
$K=\infty$ (global, extended bins) 0.814 $\pm$ 0.008 0.247 $\pm$ 0.008 0.301 $\pm$ 0.002
Mask-only pretraining 0.809 $\pm$ 0.008 0.255 $\pm$ 0.008 0.297 $\pm$ 0.002
Contrast-only pretraining 0.811 $\pm$ 0.008 0.250 $\pm$ 0.008 0.299 $\pm$ 0.002
No pretraining (random init) 0.772 $\pm$ 0.012 0.312 $\pm$ 0.014 0.268 $\pm$ 0.004
Table 6: Ablations (5 seeds). Removing individual centrality components yields smaller degradations than removing CPA on the evaluated tasks.
4.6 Data, Splits, and Overlap Removal
MoleculeNet tasks use 80/10/10 Bemis–Murcko scaffold splits (scaffolds sorted by frequency, round-robin assignment, matching DeepChem protocol; exact indices released) [8]. OGB and TDC use official splits [4, 5]. We remove any benchmark molecules from the pretraining corpus via exact match on canonical isomeric SMILES generated by our deterministic RDKit standardization pipeline (<5<5k molecules affected).
5 Discussion
Tables 35 and 1 show consistent improvements attributable to the CPA channel (objective-matched comparison to SparseGraphormer-K3) and dual pretraining (SparseGraphormer-K3 vs GraphMAE-K3). Table 6 indicates the CPA contribution is substantially larger than individual static centrality components, with degree-normalized CPA consistent with a dynamic cardinality interpretation. For typical drug-like molecules (median $N\approx 22$), $K=3$ covers nearly all nodes (median coverage $95\%$), suggesting gains arise primarily from structured biases and CPA rather than strict sparsity enforcement; on larger graphs ($\approx$15%, computed from the pretraining corpus after filtering), $K$ provides meaningful compute savings (Table 2).
6 Limitations
$K=3$ may miss long-range effects in very large molecules. Gains shown on public benchmarks; prospective validation remains future work. Our mask-based sparsity does not exploit dedicated block-sparse kernels, limiting efficiency gains to effective sequence length reduction.
7 Conclusion
CardinalGraphFormer advances structured sparse graph transformers with a simple yet effective cardinality-preserving channel, yielding consistent gains across diverse molecular benchmarks.
Code and Data Availability
Compute Resources
[…]
Appendix A Additional Results
Available in repository (appendix_results.pdf).