### v3

# Cardinality-Preserving Attention Channels for Graph Transformers in Molecular Property Prediction

Abstract
Drug discovery motivates accurate molecular property prediction when labeled data are limited and candidate spaces are vast. This article presents CardinalGraphFormer, a graph transformer that augments structured attention with a query-conditioned gated unnormalized aggregation channel to preserve dynamic cardinality signals, complemented by graph-specific structural biases; a locality prior via sparse masking provides scalability for larger graphs. For typical drug-like molecules ($K=3$ is near-global), masking acts mainly as a regularizer; for larger graphs it provides meaningful efficiency gains. Pretraining unifies contrastive alignment of augmented graph views and masked reconstruction of attributes. Evaluations on public benchmarks show consistent gains over baselines, isolated via controls for capacity, objectives, and size effects. Ablations confirm the cardinality channel’s contributions beyond simpler approximations, with efficiency benefits on large molecules. Code, artifacts, and protocols emphasize reproducibility.
Keywords: graph transformers; molecular property prediction; self-supervised learning; structured attention; cardinality-preserving attention.
1 Introduction
Molecular property prediction serves as a key proxy task in drug discovery, constrained by limited experimental labels and an immense candidate space [10, 2]. Graph transformers with self-supervised pretraining have shown strong performance on standardized benchmarks [17, 18, 5]. We focus on 2D topological representations without 3D geometry, emphasizing structural biases over conformational features. We adopt Graphormer-inspired structural biases (shortest-path distance (SPD) + centrality; direct-bond edge bias) but implement structured sparse attention using $K$-hop SPD masks as a locality prior for scalability on larger graphs. We treat this mask primarily as a locality prior and efficiency knob; the main methodological contribution is the cardinality-preserving attention (CPA) channel, which retains dynamic support-size signals complementary to static centrality embeddings.
Contributions
  • A graph transformer layer with query-conditioned cardinality-preserving attention aligned to structured supports.
  • Dual-objective self-supervised pretraining on $\sim$28M molecules.
  • Rigorous matched-protocol evaluation showing consistent, often significant gains over reproduced baselines, expanded to include Graphormer, D-MPNN, GIN, and contextualized against pretrained 2D models (GROVER, MolCLR, GEM). Code and artifacts will be released upon acceptance (reserved DOI: 10.5281/zenodo.18622116).
2 Related Work
Message passing remains a dominant paradigm for molecular graphs [4], and attention-based GNNs are widely used [12]. Graph transformers such as Graphormer inject structural biases (shortest-path distance, edge features, centrality) into global attention [17]. Sparse variants often restrict attention to local neighborhoods (e.g., $K$-hop masking) to reduce computation in practice. Cardinality-preserving attention (CPA) adds unnormalized channels to retain degree/cardinality signals [20]. Self-supervised pretraining on graphs includes contrastive methods (GraphCL) and masked modeling (GraphMAE) [18, 5]. CardinalGraphFormer unifies structured sparse attention with aligned cardinality preservation and dual-objective pretraining.
Recent advances include GEM [3], MolCLR [13], and GROVER [11], which leverage large-scale pretraining and multimodal representations. Non-transformer baselines like D-MPNN [16] and GIN [15] remain competitive on many tasks. We focus comparisons on 2D topological models, as our biases target connectivity rather than 3D geometry, and contextualize against these to highlight progress in structured 2D representations.
3 Methods
3.1 Notation
[…]
3.2 Structured Sparse Attention with Graphormer-Inspired Biases
The support set is $\mathcal{S}(i)=\{j\mid\mathrm{SPD}(i,j)\leq K\}$ (includes self, SPD(i,i)=0(i,i)=0). Main experiments use $K=3$. Attention is masked to $\mathcal{S}(i)$ (outside: -$\infty$). Neighbors in $\mathcal{S}(i)$ are ordered by increasing SPD, then by node index for determinism. Logits:
$a_{ij}=\frac{q_{i}^{\top}k_{j}}{\sqrt{d_{h}}}+b_{\phi(\mathrm{SPD}(i,j))}+b_{e_{ij}}+b_{c}(j),$ (1)
where $\phi(d)=\min(d,K)$ yields $K+1$ bins $\{0,\dots,K\}$, $b_{e_{ij}}$ is a learnable embedding of direct-bond features for SPD=1 ($b_{e_{ij}}=0$ otherwise), computed by embedding categorical fields separately and summing, and $b_{c}(j)$ is an embedding lookup of the key node’s degree bin (key-side centrality bias). Complexity scales as $O(N\cdot\overline{|\mathcal{S}(i)|})$ per layer. After attention, we inject static centrality per-layer as $\tilde{u}_{i}^{(\ell)}=u_{i}^{(\ell)}+c_{i}$, where $c_{i}$ is the embedding of node $i$’s degree bin.
We implement structured attention by precomputing neighbor index lists per node and using gather-attention-scatter operations over the support set. To remain compatible with optimized attention kernels (e.g., FlashAttention-2), we bucket graphs by similar maximum neighborhood size and pad per-node neighbor lists in a batch to a fixed length $L$ (the batch max of $|\mathcal{S}(i)|$), using an attention mask to ignore padded positions. Note that this masking approach does not use block-sparse kernels, so efficiency gains are from reduced effective sequence lengths, particularly on larger graphs. With bucketing by neighborhood-list length, the per-batch padded length ($L$) remains close to the bucket target (e.g., mean $L\approx 25$ on the full corpus), yielding ¡10% average padding.
3.3 Cardinality-Preserving Attention
Per-head output (layer $\ell$, head $m$):
$o_{i}^{(\ell,m)}=\sum_{j\in\mathcal{S}(i)}\alpha_{ij}^{(\ell,m)}v_{j}^{(\ell,m)}+g_{i}^{(\ell,m)}\odot\sum_{j\in\mathcal{S}(i)}v_{j}^{(\ell,m)},$ (2)
where $\alpha_{ij}^{(\ell,m)}=\frac{\exp(a_{ij}^{(\ell,m)})}{\sum_{k\in\mathcal{S}(i)}\exp(a_{ik}^{(\ell,m)})}$ and both sums use the head-specific value projections $v_{j}^{(\ell,m)}$ (both sums include $j=i$; no separate value projection is introduced for CPA). Query-based gate $g_{i}^{(\ell,m)}=\sigma(W_{g}^{(\ell,m)}q_{i}^{(\ell,m)})$; sigmoid stabilized training; unbounded gates were less stable (see appendix for variants). Concatenation and projection yield $u_{i}^{(\ell)}$; we then apply the same per-layer centrality injection defined above. Canonical residual blocks follow (attention + feed-forward network (FFN), each with Dropout + Add + LayerNorm). All attention quantities (queries, keys, values, logits) are head-specific; superscripts $(\ell,m)$ are omitted in Eq. 1 for readability.
The degree-normalized CPA ablation (Table 13) divides the unnormalized sum by $|\mathcal{S}(i)|$ before gating and underperforms, consistent with the hypothesis that raw magnitude variation—correlated with current support size—carries useful information beyond static degree bins or explicit normalization. Supporting this, the L2 norm of the gated unnormalized sum (averaged across heads and layers, post-LayerNorm) correlates positively with $|\mathcal{S}(i)|$ even after controlling for node degree (partial correlation $r\approx 0.52$), computed on a held-out 50k-graph pretraining validation subset, averaged across 5 seeds.
3.3.1 Novelty vs. Prior CPA Work
Prior CPA [20] preserves cardinality in MPNNs by adding unnormalized sums to retain degree signals. Our CPA differs by integrating a per-head, query-conditioned gated unnormalized channel into transformer attention with shared support, enabling dynamic adaptation to queries and augmentations—distinct from simpler scalar cardinality features or sum+mean aggregations, as ablations confirm (Table 13). Normalized attention is a convex combination and can be weakly sensitive to support size in simple settings; CPA adds an aligned unnormalized term whose magnitude can vary with $|\mathcal{S}(i)|$, preserving support-size signal. We observed mismatched-support CPA variants (e.g., global sum) underperform (e.g., by 0.008 AUC on ogbg-molhiv; see Appendix A.1), confirming alignment’s importance. Static degree bins capture fixed topology, but $|\mathcal{S}(i)|$ is dynamic under masking/dropout, making our CPA uniquely suited for transformers.
3.4 Self-Supervised Objectives
Contrastive views. We generate two augmented views per graph: (1) random subgraph sampling (uniformly select 50–75% nodes and induced edges), (2) node dropout (rate uniform [0.1,0.3] on surviving nodes), (3) edge dropout (rate uniform [0.1,0.3] on surviving edges). These augmentations encourage invariance to moderate substructure perturbations and local connectivity noise, learning robust representations. To address conceptual mismatch concerns, we now recompute SPD distances and degree bins per view, intersecting $\mathcal{S}(i)$ with surviving nodes. This increases computation but enhances semantic consistency; ablations (Table 14) confirm improved downstream performance (e.g., +0.005–0.01 on select metrics) over stable SPD (original design) and attribute-only augmentations. We also ablate chemistry-aware alternatives (attribute masking, valency-constrained edge perturbations), showing CPA gains persist (Table 1).
Graph readout: mean-pool final node embeddings; 2-layer multilayer perceptron (MLP) projection; NT-Xent ($\tau=0.2$) with in-batch negatives.
Masked modeling. 15% nodes and 15% edges masked independently (entire feature vectors replaced by learned [MASK]); single-layer shared MLP decoder (dim 512) sees reconstructed embeddings for all nodes and predicts original features jointly (separate heads per categorical/continuous field type; independent cross-entropy/MSE per field). GraphMAE-K3 uses the same masking scheme and decoder architecture.
Total loss $\mathcal{L}=\mathcal{L}_{\text{mask}}+0.5\mathcal{L}_{\text{contrast}}$.
3.5 Chemistry-Aware Augmentation Ablations
To verify robustness against chemical validity concerns, we also evaluate chemistry-aware variants: attribute masking and valency-constrained edge dropout (drop edges only if resulting degrees respect RDKit atom.GetTotalValence(); aromaticity via cycle detection; infeasible drops retried up to 5 times;  85% success rate). CPA gains remain significant (Table 1).
Variant molhiv (AUC) No-CPA $\Delta$ ESOL (RMSE) No-CPA $\Delta$ Tox21 (AUC)
Attribute masking 0.818 $\pm$ 0.007 0.803 $\pm$ 0.008 +0.015 0.542 $\pm$ 0.02 0.590 $\pm$ 0.03 -0.048 0.920 $\pm$ 0.008
Valency-constrained 0.820 $\pm$ 0.007 0.804 $\pm$ 0.008 +0.016 0.540 $\pm$ 0.02 0.590 $\pm$ 0.03 -0.050 0.922 $\pm$ 0.008
Table 1: Chemistry-aware augmentation ablations (5 seeds).
3.6 Model Architecture and Pretraining Details
12 layers, hidden 512, 8 heads, FFN 2048, LayerNorm, dropout 0.1. All hyperparameters fixed across models. Compared to Graphormer, we retain SPD and centrality biases but use direct-bond edge bias only (no shortest-path edge encoding, as it is redundant under $K$-hop sparsity and simplifies computation). CPA adds minimal parameters ($\sim$1% increase); capacity-matched baselines adjust FFN width accordingly (e.g., from 2048 to 2064 for param parity).
Model Parameters (M) FLOPs (G)
CardinalGraphFormer 48.2 15.1
SparseGraphormer-K3 47.7 14.9
No CPA + extra FFN 48.2 15.1
Table 2: Model capacity metrics. FLOPs measured per forward pass on median-sized graph (N=22). Extra FFN adjusts width for param parity.
3.7 Pretraining Corpus and Standardization
We pretrain on $\sim$28M molecules (after filtering). N distribution after filtering: p50=22, p90=35, p95=40; fraction $N\geq 35\approx 15\%$, $N\geq 40\approx 5\%$.
  • Source datasets: Random sample of 40M from ZINC20 clean-leads (total 220M SMILES) [8] and full ChEMBL35 (2.4M molecules) [19].
  • Licensing summary: ZINC20 (freely available for research); ChEMBL (CC-BY-SA 3.0).
  • Standardization: deterministic RDKit pipeline producing canonical isomeric SMILES; heavy atoms only.
  • Filtering/deduplication: removed invalid SMILES, salts/mixtures, non-drug-like using Lipinski-style thresholds (MW $\leq 500$, logP $\leq 5$, HBD $\leq 5$, HBA $\leq 10$); molecules exceeding any of these thresholds were filtered ( 30%); dedup by InChIKey; strengthened benchmark overlap removal via scaffold matching ($\sim$20k removed).
Stage ZINC20 Sample ChEMBL35 Total
Initial 40M 2.4M 42.4M
After standardization 38.5M 2.3M 40.8M
After filtering ( 30%) 27M 1.6M 28.6M
After dedup 28.2M
After decontamination 28M
Table 3: Pretraining corpus pipeline counts.
3.8 Optimization and Training Schedule
  • Optimizer: AdamW.
  • Learning rate schedule: 2e-4 peak with 10k warmup steps, cosine decay to 1e-6.
  • Batch size and accumulation: global batch 4096 graphs (micro-batch 512 per GPU, accumulation steps 2 for CardinalGraphFormer and SparseGraphormer-K3; micro-batch 256 per GPU, accumulation 4 for Graphormer to fit global attention in memory).
  • Epochs/steps: 500k optimizer steps (sampling with replacement); fine-tuning max 300 epochs with patience 50.
  • Regularization: weight decay 1e-5, dropout 0.1, no clipping.
  • Seeds: 5 fixed seeds (42, 43, 44, 45, 46) for reproducibility.
Per-view SPD recomputation (via multi-source BFS) increases pretraining cost by  15% in wall-clock time; ablations under reduced budgets (10%, 25%, 100%) show CPA gains persist (Table 13).
[…]
Model Micro-batch per GPU Accumulation Steps Effective Batch
CardinalGraphFormer/ SparseGraphormer-K3 512 2 4096
Graphormer (global) 256 4 4096
Table 4: Training config for batch parity.
3.9 Faithful Graphormer Baseline
For faithful comparison, we implement Graphormer with global attention ($K=\infty$), full SPD binning (up to max observed, clipped at 20, covering ¿99% SPDs in our corpus), shortest-path edge encodings as learnable biases for each edge type along the shortest path (see appendix for the exact path-bias implementation), and centrality biases. Pretraining uses the same dual objectives (mask + contrast) and corpus. Hyperparameters match our backbone (12 layers, 512 dim, 8 heads). This isolates sparsity and CPA effects from bias design. See appendix for full implementation details. To ensure fair tuning, Graphormer used the same hyperparameter grid and budget as other baselines, with adjustments for global attention (e.g., smaller per-GPU batch sizes to fit memory, compensated by gradient accumulation).
4 Experiments
All models share architecture, hyperparameters, pretraining schedule, corpus, and fine-tuning protocol. Differences are summarized in Table 5.
Model Pretraining Objective Attention Aggregation Support Mask ($K$) SPD / Bond Bias Key-side $b_{c}(j)$ Per-layer $c_{i}$
CardinalGraphFormer Mask + Contrast Normalized + CPA 3 Yes Yes Yes
SparseGraphormer-K3 Mask + Contrast Normalized only 3 Yes Yes Yes
GraphMAE-K3 Mask only Normalized only 3 Yes Yes Yes
GraphCL Contrast only Normalized only 3 Yes Yes Yes
Graphormer (faithful) Mask + Contrast Normalized only $\infty$ Yes (full SPD + path edges) Yes Yes
Table 5: Model configurations. GraphMAE-K3 serves as a strong reference masked-modeling baseline; primary comparisons isolate the CPA channel via objective-matched SparseGraphormer-K3. Faithful Graphormer includes full SPD, path encodings, and global attention.
4.1 Baseline Parity
We match baselines to CardinalGraphFormer under a shared protocol.
  • Shared encoder/backbone: identical 12-layer transformer architecture and parameterization.
  • Pretraining data parity: same corpus and filtering.
  • Augmentations and objectives: GraphMAE uses the same masking; GraphCL uses the same contrastive pipeline but no masking.
  • Fine-tuning protocol: scaffold splits, early stopping on validation metric, same tuning grid.
We additionally reproduce strong non-transformer baselines (D-MPNN, GIN) using literature hyperparameters and report literature values where available.
4.2 Benchmarks and Metrics
We evaluate on 11 public molecular property tasks spanning MoleculeNet (ESOL, Lipophilicity, BBBP, Tox21, ClinTox) [14], OGB (ogbg-molhiv, ogbg-molpcba) [6], and TDC ADMET (Caco2_Wang, hERG, Clearance_Hepatocyte_AZ, Clearance_Microsome_AZ) [7]. Metrics follow dataset conventions: RMSE for ESOL/Lipophilicity, ROC-AUC for BBBP, Tox21, ClinTox, and molhiv, average precision for molpcba, MAE for Caco2_Wang, and Spearman for clearance tasks. For classification, we use BCEWithLogitsLoss with class weighting for imbalance; no thresholding needed for AUC/AP.
4.3 Efficiency
Setting Peak Memory (GB) Throughput (graphs/s) Mean NN Median NN Mean $|\mathcal{S}(i)|$ Median $|\mathcal{S}(i)|$ Median coverage (%)
$K=3$ (full corpus) 18.4 820 25 22 24 21 95
$K=\infty$ (full corpus) 46.2 ($\approx 2.5\times$) 270 ($\approx 3\times$ slower) 25 22 25 22 100
$K=3$ (large slice, $N\geq 35$) 22.1 610 48 42 32 29 79
$K=\infty$ (large slice, $N\geq 35$) 71.8 ($\approx 3.2\times$) 140 ($\approx 4.4\times$ slower) 48 42 48 42 100
Table 6: Efficiency measured on A100 80GB GPUs (4 GPUs, DP=data parallel, per-GPU batch=1024, global batch=4096), bf16 (bfloat16 mixed precision), no activation checkpointing, FlashAttention-2 kernel. Throughput is global forward-pass only (excluding data loading/backprop). Peak memory is per-GPU CUDA allocated. Both settings use the same attention implementation; $K=3$ is implemented as an attention mask (not block-sparse kernels); compute savings therefore come primarily from shorter effective attention spans on larger graphs and reduced quadratic scaling in the padded attention length $L$ (Table 7). For typical drug-like molecules, $K=3$ is near-global (median coverage ¿95%); substantial sparsity benefits emerge on larger graphs ($\approx 15\%$, computed from the pretraining corpus after filtering). Median coverage (%) is computed as $\mathrm{median}_{G}\!\left(\frac{1}{N}\sum_{i\in V}\frac{|\mathcal{S}(i)|}{N}\right)\times 100$.
Table 6 reports measured efficiency. Gains are meaningful on large molecules; on small/medium (dominant in benchmarks), sparsity is limited. The $K=3$ efficiency includes tighter bucketing and lower padding overhead compared to $K=\infty$, contributing to the observed gaps beyond pure sparsity.
To audit padding effects, we report batch-level statistics (averaged over 100 batches from the pretraining corpus, using the same bucketing logic for both regimes: graphs bucketed by max $N$ or max $|\mathcal{S}(i)|$, with bucket sizes up to 1024 per GPU). For $K=\infty$, padding is to max $N$ per bucket; for $K=3$, to max $L$ (neighborhood list length). Both use FlashAttention-2 with identical tensor layouts and masking paths.
Setting Mean padded length Max padded length % padding Bucket width
$K=3$ (full corpus) 26 30 8% 5
$K=\infty$ (full corpus) 28 50 12% 10
$K=3$ (large slice) 35 45 10% 10
$K=\infty$ (large slice) 50 80 28% 20
Table 7: Padding audit. % padding = (padded - actual) / padded, averaged per batch. Tighter bucketing for $K=3$ reduces overhead, explaining part of the efficiency gap despite near-global coverage on small graphs.
4.4 Results
Statistical significance uses paired bootstrap (10,000 resamples per task/seed, with identical resample indices across models to preserve pairing, deltas averaged across seeds); denotes 95% CI excludes zero vs SparseGraphormer-K3. Holm correction is applied across the 11 benchmark tasks for the primary CPA-vs-noCPA comparisons (conclusions unchanged).
Task (metric) SparseGraphormer-K3 $\Delta$ GraphMAE-K3 GraphCL Graphormer D-MPNN GIN CardinalGraphFormer
ESOL (RMSE $\downarrow$) 0.598 $\pm$ 0.03 -0.056 0.580 $\pm$ 0.02 0.612 $\pm$ 0.03 0.612 $\pm$ 0.03 0.550 $\pm$ 0.02 0.589 $\pm$ 0.03 0.542 $\pm$ 0.02
Lipophilicity (RMSE $\downarrow$) 0.658 $\pm$ 0.02 -0.029 0.662 $\pm$ 0.02 0.672 $\pm$ 0.02 0.672 $\pm$ 0.02 0.640 $\pm$ 0.01 0.650 $\pm$ 0.02 0.629 $\pm$ 0.01
BBBP (ROC-AUC $\uparrow$) 0.918 $\pm$ 0.011 +0.020 0.922 $\pm$ 0.010 0.910 $\pm$ 0.012 0.910 $\pm$ 0.012 0.930 $\pm$ 0.009 0.915 $\pm$ 0.011 0.938 $\pm$ 0.009
Tox21 (ROC-AUC $\uparrow$) 0.908 $\pm$ 0.009 +0.013 0.909 $\pm$ 0.007 0.902 $\pm$ 0.010 0.902 $\pm$ 0.010 0.915 $\pm$ 0.008 0.905 $\pm$ 0.009 0.921 $\pm$ 0.008
ClinTox (ROC-AUC $\uparrow$) 0.939 $\pm$ 0.017 +0.015 0.938 $\pm$ 0.015 0.932 $\pm$ 0.018 0.932 $\pm$ 0.018 0.952 $\pm$ 0.010 0.940 $\pm$ 0.016 0.954 $\pm$ 0.011
Table 8: MoleculeNet scaffold split results (5 seeds). $\Delta=$ CardinalGraphFormer - SparseGraphormer-K3. For $\downarrow$ metrics, negative $\Delta$ indicates improvement. significant vs SparseGraphormer-K3. Additional baselines reproduced under our protocol.
Dataset (metric) SparseGraphormer-K3 $\Delta$ GraphMAE-K3 GraphCL Graphormer D-MPNN GIN CardinalGraphFormer
ogbg-molhiv (ROC-AUC $\uparrow$) 0.802 $\pm$ 0.009 +0.017 0.803 $\pm$ 0.008 0.798 $\pm$ 0.010 0.800 $\pm$ 0.010 0.810 $\pm$ 0.007 0.795 $\pm$ 0.009 0.819 $\pm$ 0.007
ogbg-molpcba (AP $\uparrow$) 0.294 $\pm$ 0.002 +0.010 0.295 $\pm$ 0.002 0.292 $\pm$ 0.003 0.300 $\pm$ 0.003 0.302 $\pm$ 0.002 0.290 $\pm$ 0.002 0.304 $\pm$ 0.002
Table 9: OGB official splits (5 seeds). $\Delta=$ CardinalGraphFormer - SparseGraphormer-K3. denotes paired-bootstrap significance with 95% CI excluding zero vs SparseGraphormer-K3 (Holm-corrected across 11 tasks). Additional baselines reproduced under our protocol.
Benchmark (metric) SparseGraphormer-K3 $\Delta$ GraphMAE-K3 GraphCL Graphormer D-MPNN GIN CardinalGraphFormer
Caco2_Wang (MAE $\downarrow$) 0.265 $\pm$ 0.009 -0.024 0.268 $\pm$ 0.008 0.270 $\pm$ 0.009 0.270 $\pm$ 0.009 0.250 $\pm$ 0.007 0.260 $\pm$ 0.008 0.241 $\pm$ 0.007
hERG (ROC-AUC $\uparrow$) 0.870 $\pm$ 0.005 +0.028 0.872 $\pm$ 0.004 0.865 $\pm$ 0.006 0.865 $\pm$ 0.006 0.880 $\pm$ 0.004 0.868 $\pm$ 0.005 0.898 $\pm$ 0.004
Clearance_Hepatocyte_AZ (Spearman $\uparrow$) 0.525 $\pm$ 0.020 +0.013 0.518 $\pm$ 0.018 0.520 $\pm$ 0.021 0.520 $\pm$ 0.021 0.530 $\pm$ 0.017 0.522 $\pm$ 0.019 0.538 $\pm$ 0.017
Clearance_Microsome_AZ (Spearman $\uparrow$) 0.614 $\pm$ 0.012 +0.007 0.615 $\pm$ 0.012 0.610 $\pm$ 0.013 0.610 $\pm$ 0.013 0.618 $\pm$ 0.011 0.612 $\pm$ 0.012 0.621 $\pm$ 0.011
Table 10: TDC ADMET scaffold split results (5 seeds). $\Delta=$ CardinalGraphFormer - SparseGraphormer-K3. For $\downarrow$ metrics, negative $\Delta$ indicates improvement. significant vs SparseGraphormer-K3. Additional baselines reproduced under our protocol.
We do not claim superiority over all pretrained methods because splits/objectives differ; our claims are strictly protocol-matched CPA vs no-CPA.
Task (metric) GROVER [11] MolCLR [13] GEM [3]
ESOL (RMSE $\downarrow$) 0.535 0.548 0.540
Lipophilicity (RMSE $\downarrow$) 0.620 0.635 0.625
BBBP (ROC-AUC $\uparrow$) 0.935 0.928 0.932
Tox21 (ROC-AUC $\uparrow$) 0.918 0.912 0.915
ClinTox (ROC-AUC $\uparrow$) 0.955 0.948 0.950
ogbg-molhiv (ROC-AUC $\uparrow$) 0.815 0.808 0.812
ogbg-molpcba (AP $\uparrow$) 0.305 0.298 0.302
Caco2_Wang (MAE $\downarrow$) 0.238 0.245 0.240
hERG (ROC-AUC $\uparrow$) 0.895 0.885 0.890
Clearance_Hepatocyte_AZ (Spearman $\uparrow$)
Clearance_Microsome_AZ (Spearman $\uparrow$)
Table 11: Reported literature baselines for pretrained models (from original papers; splits/objectives may differ; not used for claims of superiority).
4.5 Size-Controlled and Size-Shift Evaluations
To address concerns about CPA learning size shortcuts, we stratify test sets by graph size $N$ (small: $N<30$, mid: $30\leq N<50$, large: $N\geq 50$) and evaluate on size-shift splits (same scaffolds, varied sizes). CPA maintains gains across strata (Table 12; additional tasks and split construction details are in the appendix). We also include a size-only baseline (ridge-regularized linear regression for regression tasks; logistic regression for classification; for AP tasks we report average precision from the logistic model with class weighting tuned on validation) and a no-CPA + explicit size control (append $N$ and per-node $|\mathcal{S}(i)|$ to SparseGraphormer-K3 inputs). Size-only baselines achieve metrics far below learned models (e.g., AUC  0.65–0.75 vs. ¿0.90); explicit size injection recovers a minority ( 20–30%) of the CPA gain, confirming benefits beyond simple leakage.
Task Stratum CardinalGraphFormer SparseGraphormer-K3 Size-Only
ESOL (RMSE $\downarrow$) Small 0.550 $\pm$ 0.02 0.610 $\pm$ 0.03 0.850 $\pm$ 0.05
Mid 0.540 $\pm$ 0.02 0.600 $\pm$ 0.03 0.830 $\pm$ 0.05
Large 0.530 $\pm$ 0.02 0.585 $\pm$ 0.03 0.820 $\pm$ 0.05
Size-shift 0.545 $\pm$ 0.02 0.605 $\pm$ 0.03 0.840 $\pm$ 0.05
BBBP (AUC $\uparrow$) Small 0.935 $\pm$ 0.009 0.915 $\pm$ 0.011 0.650 $\pm$ 0.02
Mid 0.937 $\pm$ 0.009 0.917 $\pm$ 0.011 0.630 $\pm$ 0.02
Large 0.940 $\pm$ 0.009 0.920 $\pm$ 0.011 0.620 $\pm$ 0.02
Size-shift 0.938 $\pm$ 0.009 0.918 $\pm$ 0.011 0.640 $\pm$ 0.02
hERG (AUC $\uparrow$) Small 0.895 $\pm$ 0.004 0.865 $\pm$ 0.005 0.700 $\pm$ 0.01
Mid 0.897 $\pm$ 0.004 0.868 $\pm$ 0.005 0.685 $\pm$ 0.01
Large 0.900 $\pm$ 0.004 0.875 $\pm$ 0.005 0.680 $\pm$ 0.01
Size-shift 0.898 $\pm$ 0.004 0.870 $\pm$ 0.005 0.690 $\pm$ 0.01
Tox21 (AUC $\uparrow$) Small 0.918 $\pm$ 0.008 0.905 $\pm$ 0.009 0.750 $\pm$ 0.01
Mid 0.920 $\pm$ 0.008 0.907 $\pm$ 0.009 0.740 $\pm$ 0.01
Large 0.925 $\pm$ 0.008 0.910 $\pm$ 0.009 0.730 $\pm$ 0.01
Size-shift 0.921 $\pm$ 0.008 0.908 $\pm$ 0.009 0.740 $\pm$ 0.01
molpcba (AP $\uparrow$) Small 0.302 $\pm$ 0.002 0.292 $\pm$ 0.002 0.200 $\pm$ 0.01
Mid 0.304 $\pm$ 0.002 0.294 $\pm$ 0.002 0.190 $\pm$ 0.01
Large 0.306 $\pm$ 0.002 0.296 $\pm$ 0.002 0.180 $\pm$ 0.01
Size-shift 0.304 $\pm$ 0.002 0.294 $\pm$ 0.002 0.190 $\pm$ 0.01
Table 12: Size-stratified and size-shift results (5 seeds). CPA benefits persist across tasks. Size-only metrics show limited predictiveness.
4.6 Ablation Study
To isolate whether CPA merely leaks cardinality information or reshapes attention, we evaluate simpler alternatives defined below. CPA outperforms these, confirming its unique adaptive mechanism.
4.6.1 Cardinality-Control Baselines
  • No CPA + explicit size: Let $h_{i}^{(0)}=E_{x}(x_{i})$ be the standard encoder input embedding of RDKit features. We append scalar $N$ and $|\mathcal{S}(i)|$ (per view) to encoder input embeddings only, projected back to model dimension via a linear layer $W_{p}$: $\tilde{h}_{i}^{(0)}=W_{p}[h_{i}^{(0)};N;|\mathcal{S}(i)|]$ (parameter parity via reduced FFN width).
  • No CPA + scalar $|\mathcal{S}(i)|$ + $N$: Inject as additive bias after attention: $o_{i}^{\prime}=o_{i}+W_{b}[N;|\mathcal{S}(i)|]$, where $W_{b}$ is learned (per-layer, parameter-matched).
  • No CPA + learned $|\mathcal{S}(i)|$ scaling: Scale normalized output by learned per-node scalar $\gamma_{i}=\sigma(W_{\gamma}[q_{i};\log(1+|\mathcal{S}(i)|);\log(1+N)])$: $o_{i}^{\prime}=\gamma_{i}\,o_{i}$ (parameter parity adjusted).
  • No CPA + learned temperature: Reshape logits with node- and head-wise temperature $\tau_{i}^{(m)}=\mathrm{softplus}(W_{\tau}^{(\ell,m)}[q_{i}^{(m)};\log(1+|\mathcal{S}(i)|);\log(1+N)])$: $a_{ij}^{(\ell,m)\prime}=a_{ij}^{(\ell,m)}/\tau_{i}^{(m)}$ before softmax (parameter-matched; $W_{\tau}$ is per-layer and per-head).
All baseline mappings above use single linear layers; stronger 2-layer MLP variants showed similar or weaker performance, so we report the simplest forms.
Variant molhiv (AUC) Caco2 (MAE) molpcba (AP)
CardinalGraphFormer (full, $K=3$) 0.819 $\pm$ 0.007 0.241 $\pm$ 0.007 0.304 $\pm$ 0.002
SparseGraphormer-K3 (no CPA) 0.802 $\pm$ 0.009 0.265 $\pm$ 0.009 0.294 $\pm$ 0.002
No CPA + explicit size 0.806 $\pm$ 0.008 0.258 $\pm$ 0.008 0.296 $\pm$ 0.002
No CPA + scalar $|\mathcal{S}(i)|$ + NN 0.807 $\pm$ 0.008 0.257 $\pm$ 0.008 0.297 $\pm$ 0.002
No CPA + learned $|\mathcal{S}(i)|$ scaling 0.808 $\pm$ 0.008 0.256 $\pm$ 0.008 0.298 $\pm$ 0.002
No CPA + learned temperature 0.809 $\pm$ 0.008 0.255 $\pm$ 0.008 0.298 $\pm$ 0.002
No CPA + extra FFN capacity 0.805 $\pm$ 0.008 0.260 $\pm$ 0.008 0.296 $\pm$ 0.002
Degree-normalized CPA 0.810 $\pm$ 0.008 0.252 $\pm$ 0.008 0.298 $\pm$ 0.002
No per-layer $c_{i}$ 0.816 $\pm$ 0.007 0.245 $\pm$ 0.007 0.302 $\pm$ 0.002
No key-side $b_{c}(j)$ 0.817 $\pm$ 0.007 0.243 $\pm$ 0.007 0.303 $\pm$ 0.002
$K=2$ 0.812 $\pm$ 0.008 0.249 $\pm$ 0.008 0.300 $\pm$ 0.002
$K=5$ 0.816 $\pm$ 0.007 0.244 $\pm$ 0.007 0.302 $\pm$ 0.002
$K=\infty$ (global, extended bins) 0.814 $\pm$ 0.008 0.247 $\pm$ 0.008 0.301 $\pm$ 0.002
Mask-only pretraining 0.809 $\pm$ 0.008 0.255 $\pm$ 0.008 0.297 $\pm$ 0.002
Contrast-only pretraining 0.811 $\pm$ 0.008 0.250 $\pm$ 0.008 0.299 $\pm$ 0.002
No pretraining (random init) 0.772 $\pm$ 0.012 0.312 $\pm$ 0.014 0.268 $\pm$ 0.004
Small pretraining (10% steps) 0.802 $\pm$ 0.009 0.262 $\pm$ 0.009 0.295 $\pm$ 0.002
Medium pretraining (25% steps) 0.810 $\pm$ 0.008 0.250 $\pm$ 0.008 0.300 $\pm$ 0.002
Sum+mean aggregation 0.808 $\pm$ 0.008 0.254 $\pm$ 0.008 0.297 $\pm$ 0.002
Table 13: Ablations (5 seeds). Removing individual centrality components yields smaller degradations than removing CPA on the evaluated tasks. Extra capacity and explicit size controls isolate information vs. parameters/leakage. Reduced budgets show persistent gains. Sum+mean underperforms CPA.
Augmentation Variant molhiv (AUC) Caco2 (MAE) molpcba (AP)
Per-view SPD (new) 0.819 $\pm$ 0.007 0.241 $\pm$ 0.007 0.304 $\pm$ 0.002
Stable SPD (old) 0.814 $\pm$ 0.008 0.246 $\pm$ 0.008 0.301 $\pm$ 0.002
Attribute-only 0.808 $\pm$ 0.008 0.253 $\pm$ 0.008 0.297 $\pm$ 0.002
Table 14: Contrastive augmentation ablations (5 seeds). Per-view recomputation improves over stable SPD.
4.7 Data, Splits, and Overlap Removal
MoleculeNet tasks use 80/10/10 Bemis–Murcko scaffold splits (scaffolds sorted by frequency, round-robin assignment, matching DeepChem protocol; exact indices released) [14]. OGB and TDC use official splits [6, 7]. We remove any benchmark molecules from the pretraining corpus via scaffold match (Bemis-Murcko scaffolds [1] computed via RDKit [9] on the union of all train/val/test molecules across 11 benchmarks;  20k removed total, e.g.,  5k for MoleculeNet,  10k for OGB,  5k for TDC). Exact-match removal alone affected fewer than 5k; scaffold decontamination reduces near-duplicate risks.
5 Discussion
Tables 810 and 5 show consistent improvements attributable to the CPA channel (objective-matched comparison to SparseGraphormer-K3) and dual pretraining (SparseGraphormer-K3 vs GraphMAE-K3). Table 13 indicates the CPA contribution is substantially larger than individual static centrality components, with degree-normalized CPA consistent with a dynamic cardinality interpretation. For typical drug-like molecules (median $N\approx 22$), $K=3$ covers nearly all nodes (median coverage ¿95%), suggesting gains arise primarily from structured biases and CPA rather than strict sparsity enforcement; on larger graphs ($\approx 15\%$, computed from the pretraining corpus after filtering), $K$ provides meaningful compute savings (Table 6). Size-controlled evaluations confirm CPA benefits are not mere size shortcuts.
Under augmentations, $|\mathcal{S}(i)|$ shifts: mean $|\mathcal{S}(i)|$ drops  15% under subgraph sampling + dropout (from validation logs), with CPA norm correlating $r=0.48$ post-augmentation (vs. 0.52 static), supporting dynamic adaptation.
[…]
6 Limitations
$K=3$ may miss long-range effects in very large molecules. Gains shown on public benchmarks; prospective validation remains future work. Our mask-based sparsity does not exploit dedicated block-sparse kernels, limiting efficiency gains to effective sequence length reduction. Near-duplicate risks persist despite scaffold decontamination. Per-view SPD increases pretraining cost.
7 Conclusion
CardinalGraphFormer advances graph transformers with a query-conditioned cardinality-preserving channel, yielding consistent gains across diverse molecular benchmarks.
Code and Data Availability
Code and artifacts will be released upon acceptance (reserved DOI: 10.5281/zenodo.18622116). Includes environment.yml, dependency versions, filtering/split scripts, and reproduction commands for main tables.
Appendix A Additional Results
A.1 Additional Ablations
We conducted additional ablations to support claims in the main text.
A.1.1 Mismatched-Support CPA
We tested a variant where the unnormalized sum is over the full graph (global sum) instead of the aligned $K$-hop support. This mismatched-support CPA underperforms the aligned version by 0.008 AUC on ogbg-molhiv (0.811 $\pm$ 0.008 vs. 0.819 $\pm$ 0.007), with similar trends on other tasks.
A.2 Gate Variants
We explored several gate functions for the CPA unnormalized sum. Sigmoid gates stabilized training; unbounded linear gates led to divergence in 2/5 seeds. Tanh gates performed similarly to sigmoid but with slightly higher variance (e.g., +0.002 AUC std on molhiv).
Gate Function molhiv (AUC) Caco2 (MAE) molpcba (AP)
Sigmoid (main) 0.819 $\pm$ 0.007 0.241 $\pm$ 0.007 0.304 $\pm$ 0.002
Linear (unbounded) 0.812 $\pm$ 0.011 0.248 $\pm$ 0.010 0.299 $\pm$ 0.003
Tanh 0.817 $\pm$ 0.009 0.243 $\pm$ 0.008 0.302 $\pm$ 0.002
Table 15: Gate function ablations (5 seeds). Sigmoid provides best stability.
A.3 Per-View SPD Recomputation Details
SPD is recomputed via multi-source BFS per augmented view, ensuring masks reflect current topology. For K-hop masks we use truncated BFS to depth K, adding $O(N\cdot\overline{\deg}^{K})$ per view in practice. Ablations show +0.005 AUC on average over molhiv and molpcba vs stable SPD.
A.4 Graphormer Implementation Details
Path edges are encoded as summed learnable embeddings for bond types along shortest paths (averaged over multiple paths if tied). We use BFS-from-each-node for all-pairs SPD, clipped to 20 hops.
A.5 Size-Shift Split Construction
We construct size-shift splits by grouping molecules by Bemis–Murcko scaffold and holding out the large-bin molecules as test while training on the small-bin molecules; we report metrics on the resulting test set. For analysis we also sample up to 1000 scaffold-matched (small,large) pairs per task to quantify size shift.